Seawater contains approximately 3.5% salt by weight. How much seawater (in kg) contains 1 kg of salt? If the density of seawater is 1030 kg/m3, what is the volume of seawater, in liters, containing 1 kg of salt

Grade:9

1 Answers

Vikas TU
7 years ago
Let weigt of sea water be x kg.
Then,
% of Nacl (salt in sea water ) = (58.5/x) * 100 = 3.5%
solving for x we get, x = 1671.428 g or 1.671 kg.
Now for 58.5 g of salt seawater is 1671.428 g
then for 1000 g of salt seawater would be => 28571.41 g or 28.57 kg.
 
Volume of sea water = 28.57/1030 = > 0.0277 m^3 or 27.73 litre.